# A Node on the Constellation: The Role of Feminist Makerspaces in Building and Sustaining Alternative Cultures of Technology Production

### v1

@@ Abstract.
copyright: ccjournalyear: 2025doi: 10.1145/3757543conference: Computer Supported Cooperative Work; October 18-22, 2025; Bergen, Norwayisbn: 2573-0142/2025/11 - 362
@@ 1. Introduction
Over the past two decades, feminist makerspaces have emerged across the United States as vital alternatives to the dominant logics of technology production (Foster, 2019; Fox et al., 2015; Toupin, 2014). These spaces were founded in response to the exclusionary and extractive cultures of the tech industry, offering community-led infrastructures for technical practice, political organizing, and care work (Savić and Wuschitz, 2018). Often self-organized and volunteer-run, feminist makerspaces seek not only to democratize access to tools and training, but to build collective spaces where people marginalized by race, gender, and class can imagine and enact different technological futures.
@@
This paper makes three contributions to the CSCW community. First, we examine how feminist makerspaces practice long-term stewardship through distributed leadership, rotating governance, and affective labor. Second, we show how they resist the dominant logic of institutional growth, choosing instead to center solidarity with local communities as a foundation for sustainability. Third, we position these spaces as prefigurative counterinstitutions: not simply resisting exclusionary tech cultures, but actively building new ones grounded in feminist values of care, reciprocity, and collective imagination (Bardzell, 2018). This inquiry is especially urgent given the ongoing marginalization of women and people with marginalized gender identities in the tech sector (World Economic Forum Staff, 2021; O’Mara, 2022), and the increasing precarity of equity work across institutional settings. As companies retreat from DEI commitments and academic programs face growing political scrutiny (Trump, 2025; Davis, [n. d.]; Wuschitz, 2022), community-led initiatives such as feminist makerspaces remain essential yet under-resourced infrastructures of justice. Understanding how these spaces persist can illuminate broader strategies for building and sustaining counterinstitutions amid such hostile ecologies (Tandon et al., 2022).
@@ 2.1. Feminist HCI, Makerspaces and Alternative Cultures of Tech Production
Community-based spaces such as hackerspaces, makerspaces, coding boot camps, and open source technology conferences produce software, hardware, and technical knowledge that seek to disrupt the status quo (Wuschitz, 2014; Dunbar-Hester, 2020). Operating outside of institutional settings such as tech companies and universities, these spaces create alternative pathways into technological practice. For example, D’Ignazio et al. (D’ignazio and Klein, 2023) describe their transformation of a hackathon into a feminist tech festival ’Make the Breast Pump Not Suck’, which included a zine library, a baby-friendly space, a policy summit, and an art exhibition. Similarly, during the COVID-19 pandemic, Hedditch and Vyas (Hedditch and Vyas, 2021) documented an online hub where immigrants and refugee members of a feminist makerspace provided each other with social support and participated in remote making tutorials. These examples illustrate how such spaces reconfigure participation in technology by foregrounding care and accessibility.
Feminist makerspaces explicitly identify with feminist principles, offering supportive environments for women and people with marginalized gender identities to work on technical projects (Toupin, 2014). Unlike many mainstream makerspaces, they aim not just to democratize access to tools, but to challenge the norms and exclusions embedded in tech culture (Roldan et al., 2018). As Costanza-Chock argues, most makerspaces fail to disrupt systemic oppression because they assume that “providing people with tools is enough” (Costanza-Chock, 2020). In contrast, feminist makerspaces invest in learning pathways that honor the experiences and knowledge of marginalized communities. Capel et al. (Capel et al., 2020) provide an illustrative case in Australia where participants engaged in woodworking while sharing personal stories of exclusion from making practices. Themes such as housing insecurity, health challenges, and relational dynamics emerged, revealing how feminist makerspaces create spaces where storytelling and craft can intersect. In doing so, feminist makers can make historically masculine practices more accessible.
Feminist tech spaces also challenge dominant narratives of techno-heroism, individualistic, novelty-driven innovation, by emphasizing relational, community-sustaining practices (Balsamo, 2020). Fox et al. (Fox et al., 2015) frame feminism as a lens to interrogate the widespread pressure to adopt new technologies for their own sake. They document feminist makerspaces that, at times, intentionally restrict digital tools to resist this logic and revalue nondigital craft work. As Rogers (Rogers, 2016) notes, these spaces reveal a long history of marginalized communities as everyday innovators and survivalist hackers. Making in this context can include knitting circles as a radical craft, protest banner design, or circuit boards soldered from recycled materials. Here, innovation is redefined as sustaining life, not scaling products.
@@ 3.2. Interviews and Data Analysis
All interviews were conducted via a virtual meeting platform with automatic transcription software, lasting between 60 and 90 minutes. Each interview followed a protocol in which we introduced ourselves as a feminist makerspace cofounder and as researchers of feminist technologies. We told interviewees that we were conducting a study on the motivations, ideals, and activities of those who are organizing or participating in feminist makerspace activities. We opened up each interview by asking the participants to describe their background, which we intentionally left open-ended. Participants discussed their academic background, careers, ethnicity and racial identities, gender identities, and economic background. We then asked what motivated them to get involved or start a makerspace with an explicitly feminist mission, and we asked participants to describe how their involvement has changed over time. We were also interested in stories of the challenges that feminist makerspace members have faced throughout their involvement, how they addressed these challenges, and we specifically asked participants “If you could change one thing about the space, what would it be?” We also asked the interviewees to describe the values and ideals of their makerspaces and the degree to which they have seen racial justice enacted through the activities of the space.
Importantly, interviews were not differentiated based on whether the participant’s makerspace was still open at the time of the interview and we asked the same questions of all participants. The questions were open ended, allowing for customization to the status of the space, and we coded interviews by flagging relational, organizational, and infrastructural practices that lent themselves to closing or sustaining the space. We also distinguish which participants represent closed or open spaces in the findings to further distinguish between the two statuses.
@@
To analyze our findings, we draw on interpretivist perspectives that aim to translate rather than transcribe social phenomena (Huberman et al., 2014). This process centered on collaborative qualitative analysis of the interview data, with the first author initially coding each transcript, followed by an iterative process in which the coauthors reviewed and refined the codes. Through weekly meetings over the course of a year, we collectively identified emergent themes, discussed patterns across interviews, and engaged in ongoing refinement of analytic memos (Charmaz, 2006, Ch. 4). Our initial rounds of coding focused on what motivated people to start feminist makerspaces, the challenges faced by organizations to stay open, and how racial justice was reflected in the spaces’ activities. Motivation to open feminist makerspaces stemmed largely from people dealing with hostile work environments in the tech sector, where they regularly experienced sexism and racism (coded as ’racist work environment’ or ’sexist work environment’). The main challenge faced by organizations was ensuring long-term sustainability without relying on corporations and foundations (coded as ’resisting institutional reliance’ and ’community member driven.’) Supporting marginalized community members to take on leadership roles in the organization was the most common example of racial justice activities in the spaces (coded as ’centering marginalized voices.’)
@@
Importantly, we recontextualized these data through an autoethnographic lens, drawing on the first author’s reflections on co-founding and managing a feminist makerspace for over eight years. Autoethnography, as a method, enables researchers to connect personal experience with broader sociocultural and organizational dynamics, fostering a mode of inquiry that is both reflexive and situated (Ellis et al., 2011). As someone who established a nonprofit organization with an explicitly feminist mission and who has grappled with the ethical and strategic complexities of accepting corporate funding, the first author’s experience provided a critical interpretive frame for analyzing our interview data. For example, upon learning that many feminist makerspace founders did not want to accept corporate funding for their spaces, the first author wrote a reflection about how she and the cofounder of her feminist makerspace had initially accepted corporate funding and then later transitioned to a membership-driven model based on income-informed sliding scale fees. These autoethnographic reflections appear throughout our findings as a personal response to the themes that emerged from our interviews.
@@ 4.1. Bureaucratic Barriers and Fiscal Entanglements
The gentrification resulting from an influx of tech fueled a crisis of affordability for Gene’s makerspace, and therefore they relied on their personal income to sustain the space. Despite intentions to avoid reliance on major personal or corporate contributions, most spaces in our study ultimately relied on them in their early stages or times of need. Short-term investments from founders were helpful in paying the rent in the first and last months in a new space, purchasing makerspace equipment or furnishing the space. However, the long-term and recurring expenses associated with the legal entity required a broad community of committed donors that took time to establish. Such long-term, recurring expenses included equipment maintenance and repair, bookkeeping, cleaning services, materials and supplies, paying for annual tax filing, liability insurance, and directors/operators insurance. In addition to these material costs, there were also significant labor costs. Time-intensive tasks included coordinating a Board of Directors with all of its committees, meetings, and agendas. In short, starting a makerspace, like any other nonprofit organization, required upfront funding and labor. However, the long-term success of these nonprofits ultimately depended on the time investment, commitment, and resilience of a broad base of members.
@@ 4.2. Tensions of Growth and the Need to Resist Scaling
Autoethnographic reflection. During the first author’s experience in founding and maintaining a feminist makerspace for eight years, she had to wear many hats to conduct the necessary stewardship required to maintain the space. This unpaid labor quickly led to burnout. For example, during the first five years of running a feminist makerspace in the North East that is still open, the first author was responsible for a wide range of operational tasks such as paying bills, updating the website, coordinating open houses, maintaining equipment, taking out the trash, sweeping floors, and other similar tasks that required daily and weekly upkeep. Not only do these tasks require daily attention to detail (e.g., ”Is there toilet paper in the bathroom?” ”Is the laser cutter bed vacuumed out?” ”Is the 3D printer bed calibrated?” ”Has the rent been paid?”), it could also bring up interpersonal issues (”What do we do if someone sleeps in the space overnight?” ”What do we do if our pro-choice posters cause an argument?” ”Do we raise money for an elevator so that our building can be wheelchair accessible even though we do not own the building, because the landlord will not do it?” ”Who took the sewing machine home and did not return it?”) Each of these issues are persistent and complex tensions that exist in feminist makerspaces. In fact, in the first author’s own experience, this unpaid stewarding work led to burnout and the occasional need to take a personal leave of absence from the organization.
@@ 4.2.1. Pressure to Always “Grow, Grow, Grow”
One potential avenue for organizational sustainability and financial stability identified by many makerspace participants was through growing the membership base by offering tiered pricing based on income, such that a larger number of people could afford access to the space. In fact, all of the participants we interviewed wanted to establish a broad constituent base that could sustain the space through a collective fundraising approach. Additionally, small grants from foundations or government might be part of this collective approach, but many participants wanted grants to be supplementary revenue rather than having the space rely on these sources.
@@
In a Bay Area makerspace that is still open, founding members wanted a creative, shared workspace and low-income residents in the surrounding neighborhood wanted affordable access to its programs and equipment. According to Tye, a cofounder of this space:
@@
Autoethnographic reflection. In the first author’s experience of founding a feminist makerspace, a large grant was offered from a technology company in our first two years of operation. Because we needed the money and did not yet have decision-making processes in place, we decided to take the money. Although this money allowed the space to be expanded, purchase equipment, and hire a staff person, it alienated some members who initially saw the makerspace as an anticapitalist and collective effort to reimagine tech. Women, people with marginalized gender identities and other marginalized people had become involved in the space as a way to decompress and process their stressful and often sexist and racist workplaces. These workplaces were often large tech companies, financial institutions, and educational institutions which, at the same time as underpaying and underpromoting women and people of color, were funding community-based initiatives to support women and people of color. For some members, accepting funds from these institutions was seen as problematic as such efforts were viewed as publicity campaigns that would not serve to change discriminatory hiring and promotional processes. As our organization grew and developed, we stopped accepting corporate grants and instead focused on growing our membership base, and eventually it became our main source of revenue, providing enough to pay rent and part-time staff.
@@ 4.3. Deepening Relations with Neighbors and Reclaiming Making
In a shared building where there were multiple social justice organizations working alongside each other as tenants, collaborating with neighbors was often seen as the most direct way to create city-wide partnerships related to gender and racial justice. A Bay Area makerspace was located in a high-poverty area and next to a women’s shelter. As Elliott further recalled about the neighbors of this space:
@@
Because of Dana’s lived experience as a low-income, Black, trans woman, who was also in code school, she was in a ideal position to recruit more members with similar lived experience. Relatedly, Elliott realized that forging connections with their neighbors who were part of racial justice organizations led by POC would be a straightforward and sustainable approach to attracting members. However, not all valued Dana’s approach of engaging with neighbors in one-on-one conversations. Elliott continued on to share another founding member’s pushback to Dana’s community engagement:
@@
Dana thought that forging connections with their neighbors would be a straightforward and sustainable approach to attracting members. But one of the founding members, who was a white cisgender woman, disagreed with this approach because she saw it as ‘unprofessional.’ Instead, this founding member wanted to hire outside consultants to give their advice about how the feminist makerspace could be more inclusive to marginalized communities.
@@
Rather than creating a space that provided equal access to anyone, many of the makerspace participants we interviewed discussed the importance of creating a space with an equity approach to access. Centering the experiences of women, people with marginalized gender identities, and people of color is one way to decenter dominant, white, and heteronormative ways of producing technology. This emerged as a common theme across our interviews: many feminist makerspace participants explicitly welcomed discussion and interactive workshops around embodied ways of knowing. Instead of creating a race-neutral or gender-neutral space, feminist makerspace participants wanted space to explicitly address and explore the ways in which identities are intertwined with the world around us. This intersectional approach to understanding identity often challenged the gender binary and created a supportive space for transgender members. According to Zoe, a co-founder of a Pacific North West space that is now closed:

### v2

@@ Abstract.
copyright: ccjournal: PACMHCIjournalyear: 2025journalvolume: 9journalnumber: 7article: CSCW362publicationmonth: 11doi: 10.1145/3757543
@@ 1. Introduction
Over the past two decades, feminist makerspaces have emerged across the United States as vital alternatives to the dominant logics of technology production (Foster, 2019; Fox et al., 2015; Toupin, 2014). These spaces were founded in response to the exclusionary and extractive cultures of the tech industry, offering community-led infrastructures for technical practice, political organizing, and care work (Savić and Wuschitz, 2018). Often self-organized and volunteer-run, feminist makerspaces seek not only to democratize access to tools and training, but to build collective spaces where people marginalized by racism, sexism, and classism can imagine and enact different technological futures.
@@
This paper makes three contributions to the CSCW community. First, we examine how feminist makerspaces practice long-term stewardship through distributed leadership, rotating governance, and affective labor. Second, we show how they resist the dominant logic of institutional growth, choosing instead to center solidarity with local communities as a foundation for sustainability. Third, we position these spaces as prefigurative counterinstitutions: not simply resisting exclusionary tech cultures, but actively building new ones grounded in feminist values of care, reciprocity, and collective imagination (Bardzell, 2018). This inquiry is especially urgent given the ongoing marginalization of women and people of color in the tech sector (World Economic Forum Staff, 2021; O’Mara, 2022), and the increasing precarity of equity work across institutional settings. As companies retreat from DEI commitments and academic programs face growing political scrutiny (Trump, 2025; Davis, [n. d.]; Wuschitz, 2022), community-led initiatives such as feminist makerspaces remain essential yet under-resourced infrastructures of justice. Understanding how these spaces persist can illuminate broader strategies for building and sustaining counterinstitutions amid such ”hostile ecologies” (Tandon et al., 2022).
@@ 2.1. Feminist HCI, Makerspaces and Alternative Cultures of Tech Production
Community-based spaces such as hackerspaces, makerspaces, and open source technology conferences produce software, hardware, and technical knowledge that seek to disrupt the status quo (Wuschitz, 2014; Dunbar-Hester, 2020). Operating outside of institutional settings such as tech companies and universities, these spaces create alternative pathways into technological practice. For example, D’Ignazio et al. (D’ignazio and Klein, 2023) describe their transformation of a hackathon into a feminist tech festival ’Make the Breast Pump Not Suck’, which included a zine library, a baby-friendly space, a policy summit, and an art exhibition. Similarly, during the COVID-19 pandemic, Hedditch and Vyas (Hedditch and Vyas, 2021) documented an online hub where immigrants and refugee members of a feminist makerspace provided each other with social support and participated in remote making tutorials. These examples illustrate how such spaces reconfigure participation in technology by foregrounding care and accessibility.
Feminist makerspaces explicitly identify with feminist principles, offering supportive environments for women and people with marginalized gender identities to work on technical projects (Toupin, 2014). Unlike many mainstream makerspaces, they aim not just to democratize access to tools, but to challenge the norms and exclusions embedded in tech culture (Roldan et al., 2018). As Costanza-Chock argues, most makerspaces fail to disrupt systemic oppression because they assume that “providing people with tools is enough” (Costanza-Chock, 2020). In contrast, feminist makerspaces invest in learning pathways that honor the experiences and knowledge of marginalized communities. Capel et al. (Capel et al., 2020) provide an illustrative case in Australia where participants engaged in woodworking while sharing personal stories of exclusion from making practices. Themes such as housing insecurity, health challenges, and relational dynamics emerged, revealing how feminist makerspaces create spaces where storytelling and craft can intersect. In doing so, participants can make historically masculine practices more accessible.
Feminist tech spaces also challenge dominant narratives of techno-heroism, individualistic, novelty-driven innovation, by emphasizing relational, community-sustaining practices (Balsamo, 2020). Fox et al. (Fox et al., 2015) frame feminism as a lens to interrogate the widespread pressure to adopt new technologies for their own sake. They document how the activities supported by feminist makerspaces resist this logic and revalue nondigital craft work. As Rogers (Rogers, 2016) notes, these spaces reveal a long history of marginalized communities as everyday innovators and survivalist hackers. Making in this context can include knitting circles as a radical craft, protest banner design, or circuit boards soldered from recycled materials. Here, innovation is redefined as sustaining life, not scaling products.
@@ 3.2. Interviews and Data Analysis
All interviews were conducted via a virtual meeting platform with automatic transcription software, lasting between 60 and 90 minutes. Each interview followed a protocol in which we introduced ourselves as a feminist makerspace cofounder and as researchers of feminist technologies. We told interviewees that we were conducting a study on the motivations, ideals, and activities of those who are organizing or participating in feminist makerspaces. We opened up each interview by asking the participants to describe their background, which we intentionally left open-ended. Participants discussed their academic background, careers, ethnicity and racial identities, gender identities, and economic background. We then asked what motivated them to get involved or start a makerspace with an explicitly feminist mission, and we asked participants to describe how their involvement has changed over time. We were also interested in stories of the challenges that feminist makerspace members have faced throughout their involvement, how they addressed these challenges, and we specifically asked participants “If you could change one thing about the space, what would it be?” We also asked interviewees to describe the values and ideals of their makerspaces and the degree to which they have seen racial justice enacted through the activities of the space. The interviews were not differentiated based on whether the participant’s makerspace was still open at the time of the interview, and we asked the same questions of all the participants.
@@
To analyze our findings, we drew on interpretivist perspectives that aim to translate rather than transcribe social phenomena (Huberman et al., 2014). This process centered on collaborative qualitative analysis of the interview data, with the first author initially coding each transcript, followed by an iterative process in which the coauthors reviewed and refined the codes. Through weekly meetings over the course of a year, we collectively identified emergent themes, discussed patterns across interviews, and engaged in ongoing refinement of analytic memos (Charmaz, 2006). Our initial rounds of coding focused on what motivated people to start feminist makerspaces, the challenges faced by organizations to stay open, and how racial justice was reflected in the spaces’ activities. Motivation to open feminist makerspaces stemmed largely from people experiencing hostile work environments in the tech sector, where they regularly faced sexism and racism (coded as ’racist work environment’ or ’sexist work environment’). The main challenge faced by organizations was ensuring long-term sustainability without relying on corporations and foundations (coded as ’resisting institutional reliance’ and ’community member driven’). Supporting marginalized community members to take on leadership roles in the organization was the most common example of racial justice activities in the spaces (coded as ’centering marginalized voices’).
@@
Importantly, we recontextualized these data through an autoethnographic lens, drawing on the first author’s reflections on cofounding and managing a feminist makerspace for over eight years. Autoethnography, as a method, enables researchers to connect personal experience with broader sociocultural and organizational dynamics, fostering a mode of inquiry that is both reflexive and situated (Ellis et al., 2011). As someone who established a nonprofit organization with an explicitly feminist mission and who has grappled with the ethical and strategic complexities of accepting corporate funding, the first author’s experience provided a critical interpretive frame for analyzing our interview data. For example, upon learning that many feminist makerspace founders struggled with corporate and philanthropic entanglement, the first author wrote a reflection about how she and the cofounder of her feminist makerspace had initially accepted corporate funding and then later transitioned to a membership-driven model based on income-informed sliding scale fees. These autoethnographic reflections appear throughout our findings as a personal response to the themes that emerged from our interviews.
@@ 4.1. Bureaucratic Barriers and Fiscal Entanglements
The gentrification resulting from an influx of tech fueled a crisis of affordability for Gene’s makerspace, and therefore they relied on their personal income to sustain the space. Despite intentions to avoid reliance on major personal or corporate contributions, most spaces in our study ultimately relied on them in their early stages or times of need. Short-term investments from founders were helpful in paying the first month’s rent and deposit for a new space, purchasing makerspace equipment or furnishing the space. However, the long-term and recurring expenses associated with the legal entity required a broad community of committed donors that took time to establish. Such long-term, recurring expenses included equipment maintenance and repair, bookkeeping, cleaning services, materials and supplies, paying for annual tax filing, liability insurance, and directors/operators insurance. In addition to these material costs, there were also significant labor costs. Time-intensive tasks included coordinating a Board of Directors with all of its committees, meetings, and agendas. In short, starting a makerspace, like any other nonprofit organization, required upfront funding and labor. However, the long-term success of these nonprofits ultimately depended on the time investment, commitment, and resilience of a broad base of members.
@@ 4.2. Tensions of Growth and the Need to Resist Scaling
Autoethnographic reflection. During the first author’s experience in founding and maintaining a feminist makerspace for eight years, she had to wear many hats to conduct the necessary stewardship required to maintain the space. This unpaid labor quickly led to burnout. For example, during the first five years of running a feminist makerspace in the North East that is still open, the first author was responsible for a wide range of operational tasks such as paying bills, updating the website, coordinating open houses, maintaining equipment, taking out the trash, sweeping floors, and other similar tasks that required daily and weekly upkeep. Not only do these tasks require daily attention to detail (e.g., ”Is there toilet paper in the bathroom?” ”Is the laser cutter bed vacuumed out?” ”Is the 3D printer bed calibrated?” ”Has the rent been paid?”), but they can also bring up organizational issues (”What do we do if someone sleeps in the space overnight?” ”What do we do if our pro-choice posters cause an argument?” ”Do we raise money for an elevator so that our building can be wheelchair accessible even though we do not own the building, because the landlord will not do it?” ”Who took the sewing machine home and did not return it?”) Each of these issues are persistent and complex tensions that exist in feminist makerspaces. In fact, in the first author’s own experience, this unpaid stewarding work led to the occasional need to take a personal leave of absence from the organization.
@@ 4.2.1. Pressure to Always “Grow, Grow, Grow”
One potential avenue for organizational sustainability and financial stability identified by many makerspace participants was through growing the membership base by offering tiered pricing based on income, such that a larger number of people could afford access to the space. In fact, all of the participants we interviewed wanted to establish a broad constituent base that could sustain the space through a collective fundraising approach. Additionally, small grants from foundations or government might be part of this collective approach, but many participants wanted grants to be supplementary revenue rather than having the space reliant on these sources.
@@
Relatedly, in a Bay Area makerspace that is still open, founding members wanted a creative, shared workspace and low-income residents in the surrounding neighborhood wanted affordable access to its programs and equipment. According to Tye, a cofounder of this space:
@@
Autoethnographic reflection. In the first author’s experience of founding a feminist makerspace, a large grant was offered from a technology company in our first two years of operation. Because we needed the money and did not yet have collective decision-making processes in place, we decided to take the funding. Although this money allowed the space to be expanded, purchase equipment, and hire a staff person, it alienated some members who initially saw the makerspace as an anticapitalist and collective effort to reimagine tech. Women, people with marginalized gender identities and people of color had become involved in the space as a way to decompress and process their stressful and often sexist and racist workplaces. These workplaces were often large tech companies, financial institutions, and educational institutions which, at the same time as underpaying and underpromoting women and people of color, were funding community-based initiatives to support women and people of color. For some members, accepting funds from these institutions was seen as problematic as such efforts were viewed as publicity campaigns that would not serve to change discriminatory hiring and promotional processes. As our organization grew and developed, we stopped accepting corporate grants and instead focused on growing our membership base, and eventually it became our main source of revenue, providing enough to pay rent and part-time staff.
@@ 4.3. Deepening Relations with Neighbors and Reclaiming Making
In a shared building where there were multiple social justice organizations working alongside each other as tenants, collaborating with neighbors was often seen as the most direct way to create city-wide partnerships related to gender and racial justice. This Bay Area makerspace was located in a high-poverty area and next to a women’s shelter. As Elliott further recalled about one member of this space:
@@
Because of Dana’s lived experience as a low-income, Black, trans woman, who was also in code school, she was in a ideal position to recruit members with similar lived experience. Relatedly, Elliott realized that forging connections with their neighbors who were part of racial justice organizations led by POC would be a straightforward and sustainable approach to attracting members. However, not all valued Dana’s approach of engaging with neighbors in one-on-one conversations. Elliott continued on to share another founding member’s pushback to Dana’s community engagement:
@@
Dana thought that forging connections with their neighbors would be a straightforward and sustainable approach to attracting members. But one of the founding members disagreed with this approach because she saw it as ‘unprofessional.’ Instead, this founding member wanted to hire outside consultants to give their advice about how the feminist makerspace could be more inclusive to marginalized communities.
@@
Rather than creating a space that provided equal access to anyone, many of the makerspace participants we interviewed discussed the importance of creating a space with an equity approach to access. Centering the experiences of women, people with marginalized gender identities, and people of color is one way to decenter dominant, white, and heteronormative ways of producing technology. This emerged as a common theme across our interviews: many feminist makerspace participants explicitly welcomed discussion and interactive workshops around embodied ways of knowing. Instead of creating a race-neutral or gender-neutral space, feminist makerspace participants wanted space to explicitly address and explore the ways in which identities are intertwined with the world around us. This intersectional approach to understanding identity often challenged the gender binary and created a supportive space for transgender members. According to Zoe, a cofounder of a Pacific North West space that is now closed: